# PersonalAI: A Systematic Comparison of Knowledge Graph Storage and Retrieval Approaches for Personalized LLM agents

### v2

@@ 4 Experiment Set-Up
For graph construction (Memorize pipeline) and information retrieval (QA pipeline), we evaluate a series of 7B/8B and large-scale ($\gg$ 14B) language models to assess their performance in handling these tasks. The selected models include Qwen2.5 7B, DeepSeek R1 7B, Llama3.1 8B, GPT4o-min, and DeepSeek V3. To generate vector representations (embeddings) of natural language text data, we employ the multilingual E5-large model [^1] https://huggingface.co/intfloat/multilingual-e5-small.
Retrieval Algorithms
@@
Retrieval algorithms evaluation is conducted for A*, WaterCIrcles (WC) and BeamSeach (BS) as well as their combinations: ”WC + BS”, ”A* + BS”, ”A* + WC”. The values of hyperparameters for base algorithms were fixed (see Appendix F)
Graph search restrictions
@@
Additionally, we systematically varied the values of the hyperparameter governing the traversal constraints applied to the knowledge graph during algorithm execution. These constraints determine which node types are excluded from visitation: “E” prohibits access to episodic nodes, “T” restricts traversal of thesis nodes, and “O” bars visitation of object nodes. The “All” configuration imposes no restrictions, permitting unrestricted traversal across all node types.
Configurations summary
@@
Each QA configuration was evaluated on 100 question-answer pairs from the DiaASQ, HotpotQA, and TriviaQA datasets. The same large LLM was used both for generating responses within a given QA configuration and for constructing the corresponding knowledge graph to execute the QA pipeline. Consequently, for each fixed dataset/model pairing, 22 distinct QA configurations were derived. In total, 308 QA configurations were evaluated (the QA pipeline was not executed on TriviaQA/GPT4o-mini configurations due to resource constraints).
5 Evaluation
@@
Traditional statistical evaluation metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), and Meteor Universal (Denkowski and Lavie 2014) struggle to distinguish syntactically similar but semantically distinct texts. While semantic methods like BERTScore (Zhang et al. 2019) were introduced to address these limitations, our experiments reveal that BERTScore lacks sufficient differentiability, often failing to capture nuanced distinctions between correct and incorrect answers. Therefore, we adopt the LLM as a judge (Zheng et al. 2023) framework and choose Qwen2.5 7B. The judge evaluates QA pairs using a structured prompt containing the question, ground truth and model answer. It labels $1$ for correct answers and 0 for incorrect ones, and we use accuracy as our main metric. Corresponding LLM-prompts and details are provided in Appendix D.
@@
Also, for comparison of proposed QA pipeline with existing RAG and GrapRAG methods the ExactMatch metric is calculated with ignore_case and ignore_punctuation hyperparameters set to True.
6 Experiments and Results
@@
Based on the experimental results, a comparative table summarizing the best-performing QA configurations was compiled (see Table 1).
@@
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o–mini DeepSeek V3
@@
Table 1: Best QA configurations ranked by the LLM-as-a-Judge metric across all experiments. The corresponding quality score table presents the retrieval method and the type of restriction applied to the knowledge graph during traversal. Shortcuts for retrieval methods: BA – BeamSearch; AS – A*; BS + AS – hybrid of BeamSearch and A*; BS + WC – hybrid of BeamSearch and WaterCircles. Shortcuts for graph restrictions: all – no restrictions applied; E – episodic nodes excluded from traversal; T – thesis nodes excluded; O – object nodes excluded.
@@
As shown in Table 1, the results indicate that Qwen2.5 achieved the best performance (0.27) among the 7B models. Among all configurations tested, the highest overall effectiveness (0.77) was reached by setups incorporating GPT-4o-mini. Notably, top-performing 7B configurations predominantly relied on BeamSearch, especially under constraints that restricted traversal through episodic nodes. In contrast, the best DeepSeek V3 configurations frequently adopted a hybrid strategy combining BeamSearch and WaterCircles. Across high-performing configurations more broadly, BeamSearch consistently appeared as a key component within the retrieval pipeline.
@@
To evaluate the effect of the imposed constraints on the quality of the QA pipeline, we constructed two distinct distributions of values—each corresponding to a specific model/dataset pair and search algorithm. These distributions were then averaged across datasets and models for configurations exhibiting the lowest and highest JudgeScore values, as detailed in Table 2.
Restrictions
on nodes–types
7B 8B Largesize
worse–configs best–configs worse–configs best–configs worse–configs best–configs
E 0.03 0.44 0.09 0.45 0.27 0.07
T 0.84 0.12 0.64 0.31 0.20 0.73
O 0.13 0.44 0.27 0.25 0.53 0.20
@@
Table 2: Impact of various constraints, imposed during knowledge graph traversal, on the quality of QA pipeline: “worse–configs” – distribution for low quality configurations; “best–configs” – distribution for high quality configurations
@@
As demonstrated in Table 2, for 7B/8B models, the majority ($\approx$74%) of configurations yielding the lowest response quality impose restrictions on traversing thesis-type vertices. Conversely, a significant proportion of high-quality configurations restrict traversal of episodic and object vertices ($\approx$44% and $\approx$34% respectively). This suggests that thesis-type memories contain critical information for generating relevant responses, whereas the inclusion of episodic and simple memories introduces noise into the context, thereby degrading output quality. For larger-scale models, the trend differs: 53% of low-quality configurations restrict traversal of object vertices, while 73% of high-quality configurations restrict traversal of thesis-type nodes. This implies that larger models exhibit greater robustness in handling conditional generation from lengthy or noisy episodic memories, rendering thesis-based information redundant.
Retrieval algorithm 7B 8B Largesize worse–configs (w restr) best–configs (w restr) other–configs (w/o restr) worse–configs (w restr) best–configs (w restr) other–configs (w/o restr) worse–configs (w restr) best–configs (w restr) other–configs (w/o restr) WC 0.09 0.34 0.55 AS 0.1 0.175 0.14 0.29 0,36 0.41 0.23 0.36 0.33 BS 0.025 0.18 0.06 0.26 0.5 0.36 0.48 0.6 0.65 WC+BS 0.033 0.095 0.02 0.30 0.39 0.32 0.62 0.7 0.68 BS+AS 0.02 0.175 0.01 0.25 0.48 0.36 0.48 0.64 0.66 AS+WC 0.055 0.115 0.07 0.33 0.37 0.42 0.57 0.6 0.6
@@
Table 3: Stability of the implemented graph traversal algorithms when various restrictions are imposed: “worse–configs (w restr)” – low quality configurations and imposed restrictions; “best–configs (w restr)” – high quality configurations and imposed restrictions; “other–configs (w/o restr)” – configurations without restrictions on graph traversal
@@
Additionally, a comparative analysis was conducted, as presented in Table 3, to evaluate the robustness of various graph traversal algorithms. The results indicate that configurations employing 8B models in conjunction with a combined search strategy (A* and WaterCircles) demonstrate high stability, with performance degradation remaining within 4% across varying traversal constraints. In contrast, the BeamSearch algorithm exhibits high sensitivity to these constraints: suboptimal parameterization results in substantial performance loss, with JudgeScore varying by as much as 24% between optimal and non-optimal settings. However, in the case of larger-scale models, the combination of BeamSearch and WaterCircles algorithms yields more consistent performance, suggesting improved robustness at higher model capacities.
@@
A critical component of the implemented question-answering (QA) pipeline is the “NoAnswer” mechanism. This mechanism incorporates a directive in the large language model (LLM) prompt, instructing the model to output a predefined symbol if the provided context lacks sufficient information to generate a valid response. Table 4 summarizes the frequency of “NoAnswer” outputs across different model configurations, retrieval algorithms, and graph traversal constraints.
Retrieval Algorithm /
LLM-class
WC AS BS
WC +
BS
BS +
AS
AS +
WC
Restrictions
on graph search
@@
Table 4: The influence of the selected graph traversal algorithm and the imposed search restrictions on the percentage of generated “NoAnswer” stubs
@@
As demonstrated in Table 4, the analysis reveals distinct patterns in the occurrence of “NoAnswer” responses across different model configurations. For 7B models, the lowest frequency of “NoAnswer” responses is observed when employing a combined A* and WaterCircles algorithm with restricted traversal of thesis nodes. In contrast, 8B models exhibit minimal “NoAnswer” instances when utilizing the BeamSearch algorithm alongside a prohibition on episodic node traversal. For larger-scale models, the optimal performance, measured by the fewest “NoAnswer” responses, is achieved through a combined BeamSearch and WaterCircles approach without graph traversal constraints. These findings suggest that the aforementioned algorithms are more effective at extracting relevant information compared to alternative methods under the specified conditions.
@@
In summary, our framework demonstrates that the accuracy of results can be optimized by configuring the knowledge graph ontology and retrieval methods according to the available LLM and the selected QA task. A comparative analysis of existing RAG and GraphRAG methods against our proposed approach, conducted on the TriviaQA and HotpotQA datasets, is provided in Appendix H. Characteristics of constructed knowledge graphs can be observed in Appendix G.
7 Conclusion
@@
This work introduces a flexible and extensible framework for integrating structured memory into language model agents via knowledge graphs. By extending the AriGraph architecture with support for triplet, episodic and propositional nodes, we enable rich temporal and semantic representations that go beyond traditional RAG pipelines. Our system supports multiple graph traversal algorithms, including A*, WaterCircles, BeamSearch, and hybrid combinations that can be dynamically adapted to the model’s scale and task requirements. Through extensive evaluation on three benchmarks—DiaASQ, HotpotQA, and TriviaQA, we demonstrate that performance varies systematically with the choice of retrieval strategy and graph traversal constraints. For smaller-scale models (7B–8B), configurations that restrict episodic or object nodes and rely on BeamSearch yield the highest accuracy, while for larger models, hybrid methods combining BeamSearch and WaterCircles offer improved stability and robustness. Importantly, we show that thesis nodes often encode critical information, and excluding them typically degrades performance, especially in 7B models. Compared to existing RAG and GraphRAG methods, our approach demonstrates competitive or superior performance, particularly in handling temporally complex and contradictory information. Ablation studies further reveal that hybrid traversal strategies reduce sensitivity to graph constraints and lower the frequency of invalid responses. Overall, our system provides a principled architecture for long-term, structured memory in LLM agents, enabling personalized, context-aware reasoning at scale. It lays the groundwork for future extensions involving temporal filtering, relation-type prioritization, and more fine-grained memory control—all essential components for advancing robust, personalized AI systems.
8 Future work
In future work, we propose to enhance the temporal dynamics of the knowledge graph by introducing a “memory time” parameter, which will enable fine-grained filtering of triplets based on temporal proximity and relationship types. This modification will allow the system to selectively prioritize temporally proximate data or emphasize specific relationship categories, thereby improving the precision of personalized responses.

### v3

@@ IV-B Models
For memory graph construction (Memorize pipeline) and information retrieval (QA pipeline) we evaluate a series of 7B/8B and large-scale ($\gg$ 14B) language models to assess their performance in handling these tasks. The selected models include Qwen2.5 7B, DeepSeek R1 7B, Llama3.1 8B, GPT4o-min, and DeepSeek V3. To generate vector representations (embeddings) of natural language text data, we employ a multilingual E5-large model [^1] https://huggingface.co/intfloat/multilingual-e5-small.
IV-C Graph-traversal Algorithms
@@
Graph-traversal algorithms evaluation is conducted for A*, WaterCircles (WC) and BeamSearch (BS) as well as their combinations: ”WC + BS”, ”A* + BS”, ”A* + WC”. The values of hyperparameters for base algorithms were fixed (see Appendix E).
IV-D Graph-traversal restrictions
@@
Additionally, we systematically varied the values of hyperparameters governing the traversal constraints applied to the graph during algorithm execution. These constraints determine which vertex types are excluded from traversal: ”E” prohibits traversal of episodic vertices, ”T” prohibits traversal of thesis vertices, and ”O” prohibits traversal of object vertices. The ”all” key word means that restrictions about graph vertices traversal is not applied.
IV-E Summary of experiment configurations
@@
Each QA configuration was evaluated on 100 question-answer pairs from DiaASQ, HotpotQA, and TriviaQA datasets. The same LLM was used both for generating responses within a given QA configuration and for constructing the corresponding memory graph to execute the QA pipeline. Consequently, for each fixed dataset/model pairing, 22 distinct QA configurations were derived. In total, 308 QA configurations were evaluated (the QA pipeline was not executed on TriviaQA/GPT4o-mini configurations due to resource constraints).
IV-F Memory construction setting
Our implementation of memory consists of two main parts: graph part and vector part. The graph part stores textual representations of object, thesis and episodic vertices, their properties and relationships (edges) between them. Neo4j is used as the graph database for this part of the system. The vector part of memory stores vector representations (embeddings) of elements from the graph part to measure semantic similarity of texts for QA pipeline execution. Milvus is used as a vector database for this part of the system. Our memory model also implements a caching mechanism for storing intermediate results of QA pipeline components to reduce overall time spent to process incoming questions. This component utilizes two non-relational databases: Redis and MongoDB. During our experiments cache was enabled. All used databases were hosted and run on a single machine in personal docker containers. Medium size LLM (7B/8B) were used from local hosted Ollama docker container. LLM inference during memory construction and QA pipeline processing performed on a single NVIDIA TITAN RTX 24GB.
To evaluate QA pipeline we construct 14 memry graphs based on described datasets and LLMs above. Important to note, that we disable stage there we searching and deleting outdated knowledge in memory because of the experiments scope. An average speed of adding text fragments per minute to the memory for a given database configuration is 1̃.35 (with an average length of a processing texts 550-650 characters). Detailed characteristics of constructed memory graphs can be observed in Appendix G.
V Evaluation
@@
Traditional statistical evaluation metrics such as BLEU [papineni2002bleu], ROUGE [lin2004rouge], and Meteor Universal [denkowski2014meteor] struggle to distinguish syntactically similar but semantically distinct texts. While semantic methods like BERTScore [zhang2019bertscore] were introduced to address these limitations, our experiments reveal that BERTScore lacks sufficient differentiability, often failing to capture nuanced distinctions between correct and incorrect answers. Therefore, we adopt the LLM as a judge [zheng2023judging] framework and choose Qwen2.5 7B. The judge evaluates QA pairs using a structured prompt containing the question, ground truth and model answer. It labels $1$ for correct answers and 0 for incorrect ones, and we use accuracy as our main metric. Corresponding LLM-prompts and details are provided in Appendix F.
@@
Also, for comparison of proposed QA pipeline with existing RAG and GrapRAG methods the ExactMatch metric is calculated with ”ignore_case” and ”ignore_punctuation” hyperparameters set to True.
VI Experiments and Results
@@
Based on experimental results, a comparative table, summarizing best-performing QA configurations by LLM-as-a-Judge metric, was compiled (see Table II).
@@
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o-mini DeepSeek V3
@@
TABLE II: Best QA configurations ranked by LLM-as-a-Judge metric across all experiments. Corresponding cells contains LLM-as-a-Judge score, used retrieval algorithm and the type of restriction applied to the graph during traversal. Shortcuts for retrieval algorithms: BA – BeamSearch; AS – A*; BS + AS – hybrid of BeamSearch and A*; BS + WC – hybrid of BeamSearch and WaterCircles. Shortcuts for graph restrictions: all – no restrictions applied; E – episodic vertices excluded from traversal; T – thesis vertices excluded; O – object vertices excluded.
@@
As shown in Table II, Qwen2.5 achieved best performance (0.27) among 7B models. Among all evaluated configurations, highest overall effectiveness (0.77) was reached by setups incorporating GPT-4o-mini. Notably, top-performing 7B configurations predominantly relied on BeamSearch, especially under constraint that traversal restricted through episodic vertices. In contrast, best DeepSeek V3 configurations frequently adopted a hybrid strategy combining BeamSearch and WaterCircles. Across high-performing configurations more broadly, BeamSearch consistently appeared as a key component within the retrieval pipeline.
@@
To evaluate the effect of imposed constraints on the quality of the QA pipeline, we construct two distinct distributions of values, each corresponding to a specific model/dataset pair and retrieval algorithm. These distributions then were averaged across datasets and models for configurations exhibiting lowest and highest LLM-as-a-Judge values, as detailed in Table III.
Vertex types restrictions 7B 8B 14B+
worse configs best configs worse configs best configs worse configs best configs
E 3% 44% 9% 45% 27% 7%
T 84% 12% 64% 31% 20% 73%
O 13% 44% 27% 25% 53% 20%
@@
TABLE III: Impact of various constraints, imposed during memory graph traversal, on the quality of QA pipeline: ”worse configs” – distribution for low quality configurations; ”best configs” – distribution for high quality configurations
@@
As demonstrated in Table III, for 7B/8B models, majority ($\approx$74%) of configurations yielding the lowest response quality impose restrictions on traversing thesis-type vertices. Conversely, a significant proportion of high-quality configurations restrict traversal of episodic and object vertices ($\approx$44% and $\approx$34% respectively). This suggests that thesis-type memories contain critical information for generating relevant responses, whereas the inclusion of episodic and object memories introduces noise into the context, thereby degrading output quality. For larger-scale models, the trend differs: 53% of low-quality configurations restrict traversal of object vertices, while 73% of high-quality configurations restrict traversal of thesis-type vertices. This implies that larger models exhibit greater robustness in handling conditional generation from lengthy or noisy episodic memories, rendering thesis-based information redundant.
Retrieval algorithm 7B 8B 14B+ worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) WC 0.09 0.34 0.55 AS 0.1 0.175 0.14 0.29 0,36 0.41 0.23 0.36 0.33 BS 0.025 0.18 0.06 0.26 0.5 0.36 0.48 0.6 0.65 WC+BS 0.033 0.095 0.02 0.30 0.39 0.32 0.62 0.7 0.68 BS+AS 0.02 0.175 0.01 0.25 0.48 0.36 0.48 0.64 0.66 AS+WC 0.055 0.115 0.07 0.33 0.37 0.42 0.57 0.6 0.6
@@
TABLE IV: Stability of proposed retrieval algorithms when various restrictions are imposed: ”worse–configs (w restr)” – low quality configurations and imposed restrictions; ”best–configs (w restr)” – high quality configurations and imposed restrictions; ”other–configs (w/o restr)” – configurations without restrictions on graph traversal
@@
Additionally, a comparative analysis was conducted, as presented in Table IV, to evaluate the robustness of various graph traversal algorithms. The results indicate that configurations employing 8B models in conjunction with a combined search strategy (A* and WaterCircles) demonstrate high stability, with performance degradation remaining within 4% across varying traversal constraints. In contrast, the BeamSearch algorithm exhibits high sensitivity to these constraints: suboptimal parameterization results in substantial performance loss, with LLM-as-a-Judge varying by as much as 24% between optimal and non-optimal settings. However, in the case of larger-scale models, the combination of BeamSearch and WaterCircles algorithms yields more consistent performance, suggesting improved robustness at higher model capacities.
@@
A critical component of the implemented QA pipeline is the ”NoAnswer” mechanism. This mechanism incorporates a directive in the LLM prompt, instructing the model to output a predefined symbol if provided context lacks sufficient information to generate a valid response. Table V summarizes the frequency of ”NoAnswer” outputs across different model configurations, retrieval algorithms, and graph traversal constraints.
Retrieval
algorithm /
LLM size
WC AS BS WC+BS BS+AS AS+WC
Vertex types
restrictions /
LLM size
@@
TABLE V: The influence of selected retrieval algorithms and imposed search restrictions on the percentage of generated ”NoAnswer” stubs
@@
As demonstrated in Table V, the analysis reveals distinct patterns in the occurrence of ”NoAnswer” responses across different model configurations. For 7B models, lowest frequency of ”NoAnswer” responses is observed when employing a combined A* and WaterCircles algorithm with restricted traversal of thesis vertices. In contrast, 8B models exhibit minimal ”NoAnswer” instances when utilizing the BeamSearch algorithm alongside a prohibition on episodic vertices traversal. For larger-scale models, the optimal performance, measured by the fewest ”NoAnswer” responses, is achieved through a combined BeamSearch and WaterCircles approach without graph traversal constraints. These findings suggest that the aforementioned algorithms are more effective at extracting relevant information compared to alternative methods under the specified conditions.
It is also important to note the required time to process one user question by QA pipeline with a given LLM and retrieval algorithm: see Table VI.
LLM / Retrieval algorithm QA pipeline Latency (minutes) Mean:
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
WC 0.14 0.34 0.46 0.22 0.33 0.30
AS 2.24 4.68 3.51 - 2.53 3.24
BS 5.08 7.86 5.00 8.70 6.32 6.59
TABLE VI: QA pipeline Latency (in minutes), depending on LLM and retrieval algorithm with enabled caching mechanism
Table VI shows that QA pipeline with WaterCircles retriever are significantly faster. This is because that WaterCircles does not need to use vector part of memory graph with vector representations of its elements to perform traversal. Conversely, QA pipeline with BeamSearch algorithm turned out to be slower than QA pipeline with A* algorithm. This is because A* requires embeddings of memory graph elements to construct and traverse only one path on the graph, while BeamSearch requires to construct and monitor $N$ paths/candidates to select optimal traversal. The obtained Latency of the traversal algorithms is also highly dependent on the chosen vector database then configuring vector part of memory. In our experiments Milvus was used. After completing our main experiments, we evaluated read/write performance of five databases (Milvus, OpenSearch, Weaviate, ElasticSearch, and Qdrant) and found the following: (1) Qdrant was the fastest database, and (2) Qdrant in six times faster than Milvus. Therefore, we recommend specifying/using Qdrant when configuring our memory graph to reduce average required time to process single question by our proposed QA pipeline.
@@
In summary, our framework demonstrates that the accuracy can be increased by configuring the memory graph ontology and retrieval methods according to the available LLM and selected QA task. A comparative analysis of existing RAG and GraphRAG methods against our proposed approach, conducted on the TriviaQA and HotpotQA datasets, is provided in Appendix H.
VII Conclusion
@@
This work introduces a flexible and extensible framework for integrating structured memory into language model agents based on a knowledge graph. By extending AriGraph architecture with support of object, episodic, thesis vertices and hyper-adges we enable rich temporal and semantic representations that go beyond traditional RAG pipelines. Our system supports multiple retrieval algorithms, including A*, WaterCircles, BeamSearch and hybrid combinations that can be dynamically adapted to the model’s scale and task requirements. Through extensive evaluation on three benchmarks (DiaASQ, HotpotQA, and TriviaQA), we demonstrate that performance varies systematically with the choice of retrieval strategy and graph traversal constraints. For smaller-scale models (7B–8B), configurations that restrict episodic or object vertices and rely on BeamSearch yield the highest accuracy, while for larger models, hybrid methods combining BeamSearch and WaterCircles offer improved stability and robustness. Importantly, we show that thesis vertices often encode critical information, and excluding them typically degrades performance, especially in 7B models. Compared to existing RAG and GraphRAG methods, our approach demonstrates competitive or superior performance, particularly in handling temporally complex and contradictory information. Ablation studies further reveal that hybrid traversal strategies reduce sensitivity to graph constraints and lower the frequency of invalid responses. Overall, our system provides a principled architecture for long-term, structured memory in LLM agents, enabling personalized, context-aware reasoning at scale. It lays the groundwork for future extensions, involving: temporal filtering, edge-type prioritization and more fine-grained memory control.
VIII Future work
In future work, we propose to enhance the temporal dynamics of our memory graph by introducing a ”memory time” parameter, which will enable fine-grained filtering of triples based on temporal proximity and edge types. This modification will allow the system to selectively prioritize temporally proximate data or emphasize specific relationship categories, thereby improving the precision of personalized responses. Also, one possible future direction is to explore erasure-coded and locally recoverable layouts for sharding graph and vector indices across nodes, inspired by information-theoretic distributed storage, enabling fast repair and continued operation under partial server unavailability [10.1134/S1064226920120116]. In addition, we will investigate private and verifiable retrieval protocols (PIR with result verification) so an agent can query remote memory without revealing the user’s intent and can detect incorrect or malicious responses [10.48550/arXiv.2301.11730]. These mechanisms aim to make long-term personalized memory robust, secure, and auditable at scale.